1. Richard is starting a small business, and he needs a capital of P50,000. He already has P25,000, and is considering a loan the other P25,000 from one of the lending companies. Which will be the best company to loan from if richard wants to pay the lowest interest possible? Assume that all companies use ordinary interest. Explain.




a. Company A c. Company C




Rate:6% Rate:5%




Time:120 days Time:150 days





b. Company B




Rate:10%




Time:60 days





2. Billy purchased a photocopying machine on an installment plan with a down payment of P1500 and with 6 monthly payments of P3500. Find the installment price if there is an additional P100 finance charge.




3. A commercial bank charges 9% interest for 129-day loan. If the exact interest amount is P264, how much is the principal amount?





1
Expert's answer
2021-11-09T17:01:36-0500

Ordinary Interest is based on a 360-day year instead of a 365-day year.

a.For company A;

Rate:6% , Time:120 days.

Thus, Time= Number of days of a Loan/360.

Interest=25000x0.06x(120/360)

=P500

b. Company B

Rate:10%, Time:60 days

Time= Number of days of a Loan/360.

Interest=25000x0.1x(60/360)

=P416.67

c. Company C

Rate:5% ,Time:150 days

Time= Number of days of a Loan/360.

Interest=25000x0.05x(150/360)

=P520.83

Hence company B is the best to loan with since it has the lowest interest rate.

2.

Installment price

Hire purchase = deposit + total of monthly payments.

1500 + (6x3500)

=P22500

Thus, P22500 + P100

=P22600

3.

Principal amount

P = I / (RT)

Where Interest is Interest Amount, R is Rate of Interest, and T is period.

=264/ (0.09*0.3534)

= 264/0.0318

=P8299.74.